OpenStudy (anonymous):

a chemist mixes X grams of a 20% saline solution and Y grams of a 5% saline solution to make 4.5 grams of a 10% saline solution. what is the value of X?

OpenStudy (anonymous):

x+y=4.5 0.2x+0.05y=0.1(x+y) Now you have a system, solve for x and y.

OpenStudy (anonymous):

thanks soo much